OpenStudy (dls):

If x=f(t) and y=g(t) then d^2y/dx^2?

OpenStudy (dls):

\[\LARGE \frac{dy}{dx}=\frac{g'(t)}{f'(t)}\]

OpenStudy (dls):

\[\LARGE \frac{d^2y}{dx^2}=\frac{f'g''-g'f''}{(f')^2}\] im getting this..

OpenStudy (dls):

@terenzreignz @ParthKohli

Parth (parthkohli):

Seems right

OpenStudy (dls):

But is wrong.

Parth (parthkohli):

I'd have to look further. Probably the first step is incorrect.

OpenStudy (yrelhan4):

:O

OpenStudy (zarkon):

\[\LARGE \frac{d^2y}{dx^2}=\frac{f'g''-g'f''}{(f')^2}\] you are taking the derivative w.r.t. x not t. that is why this is wrong. your first derivative is correct

OpenStudy (dls):

how will we differentiate wrt x?

OpenStudy (dls):

okay,thanks I got it! just cube in denominator :D